### v2

# High Resolution Millimeter Wave Imaging Based on FMCW Radar Systems at W-Band

Abstract
In this paper, we present a unique 2D high resolution, compact, low-cost, low-weight, and highly accurate millimeter wave imagery system capable of operating in all weather conditions. We describe millimeter wave imaging process in detail and present several novel signal processing methods with their applications. To create the array, we utilize the Synthetic Aperture Radar (SAR) concept. The imagery system presented in this work, can strongly compete with Lidar systems as the resolution limit is at the same level. Furthermore, in contrast to the Lidar systems, our imagery system can operate in heavy rain and dense fog and produce high quality images.
We use our custom-made Frequency Modulated Continuous Wave (FMCW) radar operating at W-band with 33 GHz bandwidth for data collection and present the results.
Index Terms:
FMCW radar, high resolution radar imaging, SAR.
I Introduction
Radar imaging is an important topic with many applications that has been used for decades in different areas [1, 2, 3, 4]. The goal in radar imaging is to create 2D or 3D images from the targets located in the radar’s field of view. The image is the representation of the complex reflective coefficient field of the targets. Achieving higher resolution is of high importance in the field of radar imaging. The range resolution is directly related to the bandwidth of the system. Increasing the bandwidth at low frequencies such as L, C, or X bands, that have been used extensively for radar imaging, is considered to be hard to achieve. Moreover, enhancing the number of elements to obtain higher resolution in the azimuth direction, makes the system complex and increases the physical size of the system.
Recently, millimeter, sub-terahertz, and terahertz waves have been used to perform radar imaging [11, 12, 13, 14, 15, 16, 17, 18, 5, 6, 7, 8, 9, 10]. Compared to the systems operating at lower frequencies, the advantage of the systems operating at millimeter wave, sub-terahertz, and terahertz frequencies is that they are light-weight, compact, and have higher bandwidth. Operating with higher bandwidth will allow the system to obtain higher resolution in the range direction. Also, operating at higher frequencies makes it possible to create apertures with large number of elements while they occupy small physical space. This, in turn, yields high resolution in the azimuth direction.
The imagery systems operating at millimeter wave, sub-terahertz, and terahertz frequencies have numerous applications including material characterization, high resolution localization, non-destructive testing (NDT), and near-field imaging, to mention a few.
I-A Related Work
The work presented in [11, 12, 13, 14, 15, 16, 17, 18] are based on commercially available FMCW radar [1, 2, 19, 20] chips operating at 77 GHz with 4 GHz bandwidth. High resolution and high quality images have been created. However, due to the fact that the bandwidth of the systems is limited to 4 GHz, therefore, the best range resolution that we can achieve is approximately 5 cm.
In [5], the authors have developed a SiGe-Chip-based 80 GHz Frequency Modulated Continuous Wave (FMCW) radar system with 25.6 GHz bandwidth for high resolution imaging. The work is based on a low power SiGe-MMIC system operating as a FMCW radar. A large aperture has been created synthetically and high resolution images have been generated.
Another interesting work reported in the literature [6, 7] is an imagery system operating at 300 GHz with more than 40 GHz bandwidth. The system has been used in both Synthetic Aperture Radar (SAR) and Inverse SAR (ISAR) modes and has been used to create high resolution images. In [8], a FMCW radar, operating at terahertz frequencies, has been designed to create high resolution images based on Circular SAR (CSAR) technique. In [9], a FMCW radar has been designed at 220 GHz with 8 GHz bandwidth. Then the system has been used to generate high resolution radar images.
The major issue with imagery systems operating at sub-terahertz and terahertz frequencies is low output power and high free-space path-loss which makes them only suitable for near field and very close range applications.
I-B Motivation
In this paper, we present a high resolution imagery system. The system we use in this paper, is our custom-made wide-band radar system. We have chosen the W-band with frequencies ranging from 75 GHz to 108 GHz with 33 GHz bandwidth. Compared to the systems operating at lower frequencies the W-band will allow us to design a compact and light-weight system. In addition, we can reach higher bandwidth which as a result allows to enhance the range resolution limit of the system.
However, increasing the frequency range beyond the W-band will cause issues such as lower output power and higher free-space path-loss which systems presented in [6, 7, 8, 9, 10] suffer from.
Moreover, the system presented in this paper is a custom-made FMCW radar. FMCW radars are well-known for power efficiency and lower complexity compared to the pulsed-radars.
In addition to the system design and image formation, we present several signal processing techniques such as multi-spectral-based imaging, sub-array-based processing, image de-noising, and apodization which distinguishes our work from [11, 12, 13, 14, 16, 17, 18, 5, 6, 7, 8, 9, 10]. The multi-spectral-based imaging subject paves the way to applications such as material characterization. The sub-array-based processing approach provides approximately 3 dB gain which is a remarkable achievement since the origin of the gain is the signal processing unit. We further, describe the system and the imaging process in detail. Moreover, we present image de-noising. We use the state-of-the-art de-noising techniques from the field of image processing and apply it to the reconstructed images. Since the resolution of the reconstructed images is extremely high, therefore, it is plausible to bring sophisticated methods from the filed of image processing in to the field of radar imaging. We, further, utilize apodization technique. The apodization method is performed to reduce the side-lobe levels while preserving the resolution limit of the imagery system [21].
Another major difference between the work we present in this paper and the material reported in [11, 12, 13, 14, 15, 16, 17, 18, 5, 6, 7, 8, 9, 10] is related to the fact that we have collected data from multiple realistic scenarios in different weather conditions. The main purpose is to show the capability of the imagery system presented in this work in producing high resolution images from complex realistic scenarios in different weather conditions such as day, night, rain, and fog.
I-C Organization
The paper has been organized as follows. In Section II, we present the system model along with the imaging algorithm. Section III has been dedicated to the side-lobe reduction using apodization-based methods. In Section IV, we describe the multi-spectral imaging technique. Section V discusses the image de-noising process. In Section VI, we present the sub-array-based processing technique. Finally, section VII has been dedicated to the experimental results followed by the concluding remarks.
II System Model
Fig. 1 illustrates the geometry of the model we are considering in this paper. The array is created synthetically. The imagery system is a single element TX-RX FMCW radar which is moved along the horizontal direction to create a large array based on the SAR concept.
Figure 1: The geometry of the model.
In FMCW radar systems, the signal transmitted by the transmitter and received at the location of the receiver, after being reflected from the $l^{\text{th}}$ point target, and at the output of the mixer, is described as,
$s^{(l)}(t,\eta)=\sigma_{l}\;u\left(\frac{t-\tau_{l}(\eta)}{T}\right)\;e^{\displaystyle j4\pi(f_{c}+0.5\beta t)R_{l}(\eta)/c},$ (1)
where $f_{c}$ is the carrier frequency and $\sigma_{l}$ is the radar cross section for the $l^{\text{th}}$ point target. The parameter $\beta$ is given as ${b/T}$, where $b$ and $T$ stand for the bandwidth and the chirp time, respectively. Also, $u$ represents the envelope of the signal and is considered to be a rectangular window. Furthermore, $t$ and $\eta$ represent the fast- and slow-time parameters, respectively. Finally, $R_{l}$ is the radial distance between the target and the radar.
Upon taking Fourier transform from (1) with respect to the fast time parameter $t$, we obtain
$s^{(l)}(f,\eta)=\sigma_{l}\;{\text{sinc}\left(T(f-\frac{2R_{l}(\eta)}{c}\beta)\right)}e^{\displaystyle j4\pi f_{c}R_{l}(\eta)/c}.$ (2)
This step results in the localization of the energy of the signal in the range direction. The next step is to localize the energy of the signal in the azimuth direction. To accomplish this goal, we implement the back-projection algorithm [4] as follows
[…]
where $\hat{\sigma}_{l}$ is the estimated reflective coefficient of the $l^{\text{th}}$ point reflector located at the $i^{\text{th}}$ pixel.
III Apodization
The common technique for side-lobe reduction is by applying a window to the data in both range and azimuth directions. However, the side-lobe reduction based on windowing method will reduce the resolution of the system. Apodization is a method to decrease the side-lobe levels without changing the resolution limit of the system. In fact, the resolution of the system remains intact while the side-lobe levels are reduced and will be at the same level of the applied window [21].
Complex dual apodization is a more advanced apodization technique which operates on real and imaginary parts of the signal separately and generates better result compared to the apodization method [21]. In this paper, we will be using both the dual apodization and the complex dual apodization technique to reduce the side-lobe levels while preserving the resolution limit of the imaging system.
IV multi-spectral-based imaging
In this section, we present multi-spectral-based imaging. The idea is based on partitioning the whole bandwidth into 3 sub-bands. We, then, create 3 separate images and assign red, green, and blue colors to them, respectively. At the end, we display the image in color.
In fact, the sub-bands are created using the IF signal (also known as the beat signal). We divided the beat signal, in fast time domain, into three equal sections and each section represents one-third of the initial signal. Next, each part is used to create an independent image from the scene. Finally, all three images are assigned red, blue, and green colors and the result is plotted.
Figure 2: Dividing the whole bandwidth of the system into 3 equal sub-bands for multi-spectral-based imaging.
Fig. 2 shows the process of dividing the entire bandwidth of the system into 3 sub-bands. One of the applications of multi-spectral-based imaging is for materiel characterization in which targets that display the same response over the whole frequency band are presented as white and targets that have frequency-dependent response will be seen as non-white in the processed image.
V image de-noising
The image created by the system is corrupted with noise. In this section, we address the image de-noising process. The goal of image de-noising is to remove the noise while prevent image smearing and preserve the fine structures in the image. Image de-noising helps with observing the fine details of the image which have been covered by the noise of the system. Since the imagery system presented in this paper creates incredibly high resolution images, which is comparable to optical images in resolution and quality, therefore, we will be using image de-noising techniques from the filed of image processing.
[…]
In (4), $f$ represents the noisy image and $\lambda$ is the hyper-parameter of the optimization problem. Also, $\Omega$ represents the image domain.
The optimization problem given in (4) is convex [23] and can be solved using existing packages such as cvx in MATLAB or cvxpy in Python. However, for images with large number of pixels, solving the optimization problem, given in (4), is computationally prohibited.
To mitigate this issue, we will be solving the Euler-Lagrange equation [24] for the optimization problem given in (4). Moreover, to make the algorithm resilient to edge smearing, we will use an anisotropic version of (4). Consequently, upon using the Euler-Lagrange equation we obtain the following nonlinear elliptic partial differential equation,
$\begin{cases}&\displaystyle\frac{\partial u}{\partial t}=\nabla.\left({c(|\nabla(G_{\sigma}*u)|^{2})}{\nabla u}\right),\\ &u(x,0)=f(x),\\ &c(||\nabla I||)=e^{-\displaystyle{\left(||\nabla I||/K\right)}^{2}},\\ &G_{\sigma}(I)=\sigma^{-1/2}e^{\displaystyle-|I|^{2}/(4\sigma)},\\ &(t,x)\in(0,T)\times\Omega.\end{cases}$ (5)
In (LABEL:denoising_opt_EL), the parameter K controls the sensitivity to the edges and is a hyper-parameter and * represents the convolution operator. The algorithm presented in (LABEL:denoising_opt_EL) is in fact the nonlinear diffusion method, also known as modified Perona-Malik diffusion technique [25], which is used for image de-noising . We will implement (LABEL:denoising_opt_EL) directly in Python to achieve the de-noised image. One of the main features of the de-noising technique given in (LABEL:denoising_opt_EL) is the ability of the algorithm to preserve the edges and fine details of the image while remove the noise.
VI sub-array-based processing
In this section, we discuss sub-array-based processing. The goal of the sub-array-based method is to increase the Signal to Noise Ratio (SNR) which is a critical factor for image reconstruction at millimeter wave frequencies. We partition the array into smaller sub-arrays. Then, we perform beam-forming per each sub-array and direct the sub-beams toward a specific spot in the field of view of the imagery system. Next, we utilize all the sub-beams and perform the final beam-forming stage over all the sub-beams that we have created and generate the final image. Fig. 3 illustrates the idea behind the sub-array-based imaging.
Creating sub-arrays this way will increase the SNR which, in turn, will reduce the noise floor and that is the main goal for implementing the sub-array-based approach. The sub-arrays are created with non-zero overlap which forces the noise of the system to become correlated. Therefore, the gain we achieve is not exactly equal to the number of the elements in the sub-arrays. However, we can obtain approximately 3 dB gain in SNR which is a remarkable achievement. It should be noted that, the gain is purely obtained from the signal processing unit and that is what makes this approach valuable.
It is worth mentioning that, increasing the number of elements in the sub-array comes at the cost of narrowing the field of view of the imagery system.
Figure 3: Partitioning the whole array into sub-arrays and direct the sub-beams toward a specific place in the field of view of the imagery system which will be used to perform the final stage of the beam-forming process.
VII experimental results
In this section, we present our experimental results. We have designed a FMCW radar at W-band operating at frequencies ranging from 75 GHz to 108 GHz with 33 GHz bandwidth. .
Figure 4: The block diagram of the FMCW radar system.
Fig. 4 shows the block diagram of the developed broadband imaging radar system incorporating FMCW technology, based on a Phase-Locked Loop (PLL) FMCW generator. This generator outputs a signal spanning a broad initial frequency spectrum of 12.5 GHz to 18 GHz, with a power specification of 10 dBm. Subsequent frequency multiplication through a configured doubler and tripler extends the signal’s frequency reach, achieving an expansive band from 75 GHz to 108 GHz.
This frequency multiplication process enhances the bandwidth of the radar system which is critical for achieving high-resolution imaging. Amplification stages subsequent to each multiplication step ensure proper signal power at the output and maintain the broadband characteristic of the output. A power splitter, serving both the transmission arm and the local oscillator input for the mixer in the receive arm. The transmission arm leverages the broadband capabilities to radiate a signal through either an open wave-guide or a horn antenna depending on the scenario under test to support high-resolution imaging applications.
The receiving arm utilizes a Low Noise Amplifier (LNA) with 25 dB gain and 4 dB noise figure to amplify the reflected broadband signals. These signals, upon mixing with the local oscillator signal, yield an Intermediate Frequency (IF) that contains the beat frequencies from the different targets.
To assure the integrity of the IF signal, a high-pass filter is employed to discard undesirable low-frequency components, with subsequent stages of amplification and filtering enhancing the signal quality. The IF signal is then converted to digital format via an Analog-to-Digital Converter (ADC), making it amenable to perform complex digital processing algorithms.
The purpose of the high pass filter is to remove or decrease the energy of the close-by targets, that are within the minimum range of the radar, as well as blocking the leakage from TX into RX to protect the receiver from saturation and increase the dynamic range of the system.
The ADC block contains the ADC as well as the low pass Anti Aliasing Filter (AAF). The sampling frequency for the ADC is 5 MHz. The chirp time has been set to 5 msec.
Refer to captionNormalized Power [dB]Range [m](a)
Refer to captionNormalized Power [dB]Range [m](b)
Figure 5: a) the impulse response of the system using a corner reflector at 2.05 m radial distance, b) the 3 dB beam-width limit. .
In Fig. 5, we have shown the impulse response of the system using a corner reflector located at 2.05 m radial distance.
To perform the image reconstruction process, we have chosen frequencies ranging from 78 GHz to 102 GHz which is the flat part of the spectrum that the system provides and is equivalent to 24 GHz bandwidth. The 24 GHz bandwidth will provide us with 6.3 mm resolution in the range direction. The length of the array, for all the experiments that we present in this section, is 40 cm.
Fig. 6 illustrates the power spectrum of the system over the entire bandwidth ranging from 75 GHz to 108 GHz. It is clear from Fig. 6 that we have large power fluctuation in the beginning and at the end of the spectrum. For the imaging process, we have chosen the part corresponding to range of frequencies from 78 GHz to 102 GHz which is equivalent to 24 GHz bandwidth.
Refer to captionNormalized Power [dB]Frequency [GHz]
Figure 6: The power spectrum for the system over the entire bandwidth ranging from 75 GHz to 108 GHz.
Figure 7: The radar system as well as the motorized scanner to create the synthetic aperture.
Fig. 7 illustrates the radar system as well as the motorized scanner to create the synthetic aperture.
Figure 8: The experimental set-up.
Fig. 8 depicts the experimental set-up in which a bike has been located in front of the radar system.
VII-A Image Formation
In this subsection, we present the result of applying the back-projection algorithm to the data-set collected from a bike in front of the imaging system.
Refer to captionRange [m]Cross Range [m]
Figure 9: The result of applying the back-projection algorithm to the experimental data gathered from the bike.
Fig. 9 shows the result of applying the back-projection algorithm, given in (3), to the experimental data gathered from the bike.
Refer to captionRange [m]Cross Range [m]
Figure 10: The reconstructed image after applying Hamming window in both the range and cross-range directions.
Fig. 10 illustrates the reconstructed image while we have applied Hamming window in both range and cross-range directions to reduce the side-lobe levels. From Fig. 9 and Fig. 10, we can see all the details of the bike such as the saddle, tires, pedal, chain, spokes, gears, handlebars, and crossbar.
Refer to captionRange [m]Cross Range [m]
Figure 11: The point-cloud for the image shown in Fig. 9, the information about the intensity is included.
In Fig. 11, we have shown the point-cloud for the reconstructed image shown in Fig. 9. Fig. 11 includes the information about the intensity of the point reflectors
Refer to captionRange [m]Cross Range [m]
Figure 12: The point-cloud for the image shown in Fig. 9 with all points represented at the same intensity level.
In Fig. 12 we have shown all the point reflectors with the same intensity level.
VII-B Apodization
In this sub-section, the result of the dual apodization and complex dual apodization methods are presented. We have used rectangular and Hamming window to perform the apodization method. The rectangular window represents the image without windowing. It is important to mention that multiple windows can be used for apodization which create multi-layer algorithm for side-lobe reduction while preserving the resolution limit of the system.
The dual apodization method is based on the absolute value of the reconstructed image and provides reasonable result. However, the complex dual apodization works on the real and imaginary part of the reconstructed image and yields an incredibly better result.
Fig. 13 shows the result of utilizing the apodization methods. Fig. 13-(a) illustrates the image without apodization. In Fig. 13-(b), the result of dual apodization which has been applied to the image is shown. As can be seen from Fig. 13-(b), the side-lobe levels are lower and the resolution limit is intact. The side-lobe levels will be at the same level as the case in which the Hamming window is used.
Finally, Fig. 13-(c) depicts the result of applying the complex dual apodization technique. As it is clear from Fig. 13-(c), the result of the complex dual apodization is by far better than that of the dual apodization. Of course, the computational complexity of the complex dual apodization is slightly higher.
Refer to captionRange [m]Cross Range [m](a)
Refer to captionRange [m]Cross Range [m](b)
Refer to captionRange [m]Cross Range [m](c)
Figure 13: a) the resulting image without apodization, b) the reconstructed image after applying the dual apodization method, c) the reconstructed image after applying the complex dual apodization method.
VII-C Image De-noising
In this sub-section, we implement the modified Perona-Malik method, given in (LABEL:denoising_opt_EL), directly in Python to remove the noise from the reconstructed images.
The result of the de-noised image has been shown in Fig. 14. In order for the effect of the de-noising process to be seen clearly, we have shown all the images with 50 dB dynamic range.
Refer to captionRange [m]Cross Range [m](a)
Refer to captionRange [m]Cross Range [m](b)
Refer to captionRange [m]Cross Range [m](c)
Figure 14: a) the noisy image, b) the de-noised image, c) the de-noised image by applying the de-noising technique given in (LABEL:denoising_opt_EL) to the result of the complex dual apodization method shown in Fig. 13-(c).
As can be seen from Fig. 14-(b) the de-noising algorithm provides a remarkable result. The fine structures of the target can be uncovered from the noise while the edges have been preserved. The result shown in Fig. 14-(b) has been obtained by applying (LABEL:denoising_opt_EL) to the windowed image shown in Fig. 14-(a). In Fig. 14-(a), we have used Hamming window in both the range and azimuth directions.
In Fig. 14-(c), we have applied the de-noising algorithm presented in (LABEL:denoising_opt_EL) to the result of the complex dual apodization which has been shown in Fig. 13-(c). Since the complex dual apodization reduces the side-lobe levels while preserves the resolution limit of the system, therefore, Fig. 14-(c) provides better results compared to what we have achieved in Fig. 14-(b). In other words, more fine structures related to the bike can be seen in Fig. 14-(c) compared to what is visible in Fig. 14-(b).
VII-D Sub-Array-Based Processing
This subsection discusses the sub-array-based processing results. Fig. 15 shows remarkable results for implementing the sub-array-based processing.
We have partitioned the whole array into sub-arrays with length 10 and with 9 elements overlap between them. Subsequently, we use all these sub-arrays to perform the final stage of the beam-forming process and obtain the image.
In Fig. 15-(a), we have steered the beam toward left by 15 degree. In Fig. 15-(b), we are illuminating the bore-sight and in Fig. 15-(c), we have directed the beam toward the right by 15 degree.
By performing sub-array-based processing with sub-arrays of length equal to 10, we can achieve approximately 3 dB gain which is impressive. The reason that we can not obtain 10 dB gain, which is the real number for creating sub-arrays with the length equal to 10, is that when we create the sub-arrays, the overlap between them will force the noise of the system to become correlated. As a result, when we perform the final stage of the beam-forming process to generate the final image we will be deprived of the full array gain because of the correlation between the noise samples.
The results that we have presented in Fig. 15 are based on 50 dB dynamic range and in gray scale, thus, we can compare them with the reconstructed image shown in Fig. 14-(a).
The complete analysis of the gain of the sub-array-based processing and its dependency on the length of the sub-arrays, will deviate us from the main concept of this paper and will be the subject of our future work.
It should be noted that, the gain that we obtain by using the sub-array-based processing technique, which we have presented in this section, is exclusively the gain that is provided by the signal processing unit and that is the reason that the sub-array-based method is considered to be highly valuable.
However, as we mentioned before and as it is clear form the results presented in Fig. 15, when we increase the number of elements per each sub-array the overall field of view of the system becomes narrower. In other words, achieving higher gain using the sub-array-based approach comes at the cost of narrowing the field of view of the imagery system. As a result, to be able to perform imaging for the entire field of view of the imagery system, we need to steer the beam similar to what we have done in Fig. 15.
Refer to captionRange [m]Cross Range [m](a)
Refer to captionRange [m]Cross Range [m](b)
Refer to captionRange [m]Cross Range [m](c)
Figure 15: The result of sub-array-based processing. Sub-arrays of length equal to 10 have been created and all the sub-beams have been directed towards the location of the bike. Finally,all the elements of the array have been used to perform the beam-forming, a) steering the beam toward left by 15 deg, b) steering the beam toward the center c) steering the beam toward right by 15 deg.
VII-E Multi-Spectral-Based Imaging
The result of multi-spectral-based imaging based on 3 sub-bands has been shown in Fig. 16. The result is based on 3 separate images, one image per each sub-band. Next, they are assigned to red, green, and blue colors and the final image is created. For better visualization, we have presented $1-|\text{img}|$ which implies that if the target has the same response over the entire bandwidth, it will be shown in black color.
In order to increase the visibility of the image, we have performed contrast enhancement as well [26]. We have applied the de-noising algorithm given in (LABEL:denoising_opt_EL) to the image prior to partitioning the data into 3 sub-bands.
In Fig. 16, the target with darker color shows the same response over the entire bandwidth of the system. In other words, its reflective coefficient has less dependency on the range of frequencies that the system is operating over.
Refer to captionRange [m]Cross Range [m]
Figure 16: The result of the multi-spectral-based imaging. The 24 GHz bandwidth has been partitioned into 3 sub-bands.
The multi-spectral-based imaging can be used for material characterization and this will be the subject for our future work.
Refer to caption(a)
Refer to captionRange [m]Cross Range [m](b)
Figure 17: a) the experimental set-up for a car and two bikes in front of the radar system, b) the reconstructed image.
VII-F More Results
In this subsection, we present more experimental results from targets such as bike as well as car at near and far distances from the imagery system. It should be noted that, for all the experimental results presented in this subsection, we have used the back-projection algorithm, given in (3), to reconstruct the images. Also, we have used the de-noising algorithm given in (LABEL:denoising_opt_EL) to de-noise the images.
We have performed a test in which a car and two bikes have been located in front of the radar system. Fig. 17-(a) illustrates the experimental set-up. In Fig. 17-(b), we have presented the reconstructed image. From Fig. 17-(b), we can see all the details of all targets present in the radar’s field of view.
Refer to caption(a)
Refer to captionRange [m]Cross Range [m](b)
Figure 18: a) the experimental set-up for 2 cars, the data has been collected during night time, b) the reconstructed image.
We have conducted a different experiment which includes 2 cars in front of the imagery system. Fig. 18-(a) illustrates the experimental set-up. As can be seen from Fig. 18-(a) , the data has been collected during night time. In Fig. 18-(b), we have shown the reconstructed image. From Fig. 18-(b), we can see the details of both cars such as the bumper, headlights, license plate, and side mirror.
We have conducted an additional test. The data for this scenario has been collected in rainy weather. Fig. 19-(a) illustrates the experimental set-up for the bike situated on the ground in front of the car. As we mentioned, the data has been collected in a rainy morning in order to test the ability of the imagery system to create a high resolution image from the targets even in rainy condition. As can be seen from Fig. 19-(a), the ground, the bike, and the car are all wet and rain is coming during the period that the data is being collected. Fig. 19-(b) shows the reconstructed image.
Refer to caption(a)
Refer to captionRange [m]Cross Range [m](b)
Figure 19: a) the experimental set-up for the bike in front of the car, the data has been collected in a rainy morning, b) the reconstructed image.
To further demonstrate the all-weather capability of the imagery system, we have conducted a test in a foggy environment. The data for this scenario has been collected in dense fog. Fig. 20-(a) shows the experimental set-up for the bike situated on the ground in front of the imagery system in clear weather condition and Fig. 20-(b) presents the reconstructed image from the scene. The set-up is inside the garage and we have then created a thick fogy situation. In Fig. 21-(a), the test set-up in the fog has been shown. The thick layer of fog makes it impossible for the bike to be seen in the optical image presented by Fig. 21-(a). In Fig. 21-(b), we have presented the reconstructed image form the fogy test set-up. It is clear form In Fig. 21-(b) that even in thick fogy situation the imagery system is capable of producing high resolution image form the target. In fact the reconstructed images under no-fog and foggy conditions, presented in Fig. 20-(b) and Fig. 21-(b), respectively, are identical.
Refer to caption(a)
Refer to captionRange [m]Cross Range [m](b)
Figure 20: a) the experimental set-up for the bike in front of the imagery system, b) the reconstructed image.
Refer to caption(a)
Refer to captionRange [m]Cross Range [m](b)
Figure 21: a) the experimental set-up presented in Fig. 20-(a), the data has been collected in dense fog, b) the reconstructed image.
The next test experiment, which has been shown in Fig. 22-(a), includes 2 cars, 2 bikes, a ladder, and a scooter. Fig. 22-(b) illustrates the reconstructed image which clearly shows all the targets. The reason we have conducted this experiment is that the number of targets are higher and the targets have been located in close proximity of one another. Consequently, we can see some undesired effects that have been created by multi-path.
Refer to caption(a)
Refer to captionRange [m]Cross Range [m](b)
Figure 22: a) the experimental set-up for 2 cars, 2 bikes, a ladder, and a scooter, b) the reconstructed image.
Finally, we conclude this section with the experimental set-up for the bike located at 8.5 m distance in front of the imagery system. Fig. 23-(a) depicts the experimental result and in Fig. 23-(b) we have shown the reconstructed image. When we place the targets at far distances from the imagery system, the SNR as well as the resolution limit of the system in the cross range direction decrease.
It should be noted that, for this experiment we have used horn antennas on both the transmit and receive sides in order to increase the maximum range of the imagery system. Consequently, the field of view of the imagery system has been reduced and in the reconstructed image, which has been shown in Fig. 23-(b), we can only see the bike.
Refer to caption(a)
Refer to captionRange [m]Cross Range [m](b)
Figure 23: a) the bike in front of the imagery system at far distance, b) the reconstructed image.
VIII Conclusion
In this paper, we presented our custom-made FMCW radar system which operates at millimeter wave frequencies. We described the functionalities of the system in detail. We further, presented the imaging algorithm and several signal processing techniques for image formation and image post processing. Finally, we presented all the results based on the experimental data gathered from the imagery system. The paper is a proof that images with high resolution and high quality can be generated using compact, low-weight, and low-cost millimeter wave systems such as the one presented in this work.